Suppose that the economy is characterized by behavioral equations. C =c +yd


Where consumption autonomous= 280

Investment=270

Government spending =300

Taxes =20%

Marginal propensity to save= 40%


Solve for

a. Equilibrium GDP

b. Disposable income

C. Consumption spending



1
Expert's answer
2021-06-14T15:11:31-0400

a. GDP = Consumption + Investment + Government Spending + Net Exports

GDP = 280 + 270 + 300 + 0=850


b. Disposable Income = Income – Income Taxes

Disposable Income = "270- \frac{20}{100}*270=216"


c. Consumption spending = Marginal propensity to save x Disposable Income +consumption autonomous

Consumption spending "=\frac{40}{100}*216 +280 = 366.4"